OpenStudy (twizttiez):

Which is an equation in point-slope form for the given point and slope? Point: (2,-6) ; Slope:-3/4 A.y-6=-3/4(x+2) B.y-6=3/4(x-2) C.y+6=-3/4(x-2) D.y+6=-3/4(x+2)

TheSmartOne (thesmartone):

Ok \(\Huge y-y_1=m(x-x_1)\)

TheSmartOne (thesmartone):

so plug in the points \(\Huge (x_1,y_1)\) which are \(\Huge (2,-6)\)

TheSmartOne (thesmartone):

and the slope or m is \(\Huge \frac{-3}{4}\)

TheSmartOne (thesmartone):

So just plug it in.

OpenStudy (twizttiez):

Sorry didn't mean to ignore you @TheSmartOne i was busy

TheSmartOne (thesmartone):

Its okay.

TheSmartOne (thesmartone):

Did you ever plug it in? xD

OpenStudy (twizttiez):

I actually did XD